what is the equation of -7.7x + 1.7 = -62.3 - 1.3x into fractions if possible
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

what is the equation of -7.7x + 1.7 = -62.3 - 1.3x into fractions if possible
------------------
What is the equation of an equation? Is that really your question?

Answer by stanbon(75887)

what is the equation of -7.7x + 1.7 = -62.3 - 1.3x into fractions if possible
--------------
-(77/10)x + 17/10 = -623/10 - 13/10x
=========================================
Cheers,
Stan H.
=========================================